AnakKrakatau1883:

Which of the following inequalities matches the graph? graph of an inequality with a dashed line through the points 0, 7 and 1, 5 with shading below the line 2x + y < 7 2x − y > 7 2x + y < −7 The correct inequality is not listed

AnakKrakatau1883:

1 attachment
link1234:

2x − y > 7

XioGonz:

@link1234 wrote:
2x − y > 7
Guide him/her through the steps

AnakKrakatau1883:

Thanks again!

link1234:

ok?

XioGonz:

@link1234 wrote:
ok?
Direct answers are not allowed

link1234:

@xiogonz wrote:
@link1234 wrote:
2x − y > 7
Guide him/her through the steps
u do it i already anwerseed the question

XioGonz:

@link1234 wrote:
@xiogonz wrote:
@link1234 wrote:
2x − y > 7
Guide him/her through the steps
u do it i already anwerseed the question
You gave the answer-You have to guide through

link1234:

there Click here to get an answer to your question ✍️ Find three solutions of the equation: 2x + y = 7. 1 answer · Top answer: The given equation is 2x + y = 7 ... (1) Putting x = 0 in (1), we get, 2 × 0 + y = 7 x = 7/2 (7/2 , 0 ) is a solution of (1).Further, putting y = 1 in ... Missing: guided ‎| Must include: guided

XioGonz:

@umm

XioGonz:

@link1234 wrote:
there Click here to get an answer to your question ✍️ Find three solutions of the equation: 2x + y = 7. 1 answer · Top answer: The given equation is 2x + y = 7 ... (1) Putting x = 0 in (1), we get, 2 × 0 + y = 7 x = 7/2 (7/2 , 0 ) is a solution of (1).Further, putting y = 1 in ... Missing: guided ‎| Must include: guided
It must be your own explanation-Not one off another website >.>

link1234:

sorry i don't know then

link1234:

@link1234 wrote:
sorry i don't know then
i so sorry

AnakKrakatau1883:

I forgive you.

link1234:

tysm

AnakKrakatau1883:

Your welcome.

XioGonz:

@link1234 wrote:
@link1234 wrote:
sorry i don't know then
i so sorry
For future reference if you don't know the answer to a question it's best not to answer at all

XioGonz:

It can result in a warning

link1234:

sorry

link1234:

AnakKrakatau1883 i may get kicked out of question cove

link1234:

forever

AnakKrakatau1883:

I'm sorry if that happens. Respects go to you, bud.

link1234:

tysm :)

AnakKrakatau1883:

No problem!